OpenStudy (anonymous):

A wire 24inches long is to be cut into four pieces to form a rectangle whose shortest side has a length of x: Express the area A of the rectangle as a function of x

OpenStudy (anonymous):

Length=x so Breadth = 24-2x/2=12-x Area= x(12-x)=12x-x^2